# ND-SDF: Learning Normal Deflection Fields for High-Fidelity Indoor Reconstruction

### v2

@@ 1 Introduction
3D surface reconstruction Chen et al. (2024); Ye et al. (2024b; a); Wang et al. (2024); Ye et al. (2022; 2023b); Liu et al. (2021); Li et al. (2020) aims to recover watertight, dense 3D geometry from multi-view images Hartley & Zisserman (2003), representing a significant research area in computer vision and graphics. The recovered surfaces have proven invaluable for many downstream tasks, including robotic navigation, AR/VR, and smart cities.
@@
However, recovering high-fidelity surfaces remains challenging, as relying solely on color images for supervision often results in an underconstrained problem, especially for regions like textureless walls and ceilings. Previous methods Yu et al. (2022b); Wang et al. (2022) have attempted to mitigate this issue by incorporating auxiliary supervision. For instance, MonoSDF employs monocular cues from a pretrained model as pseudo-ground truth to supervise the reconstruction process, which alleviates the issue in textureless regions. However, the large errors due to domain gaps in monocular cues and the inconsistencies in view-dependent prior guidance often lead to visible detail loss and erroneous surfaces Yu et al. (2022b), as depicted in Figure 4 of the experimental section.
@@ 2 Related Work
Recently, the representation of 3D scenes using neural fields has gained popularity due to their expressiveness and simplicity. NeRF-type approaches Mildenhall et al. (2021); Barron et al. (2022); Müller et al. (2022); Fridovich-Keil et al. (2022); Sun et al. (2022) have proposed encoding coordinate-based density and appearance of scenes by utilizing simple multilayer perceptrons (MLPs) and explicit structures such as voxel grids, resulting in photorealistic novel view synthesis. However, these approaches fail to accurately recover surfaces due to the lack of constraints on density. To address this issue, subsequent works, such as VolSDF Yariv et al. (2021) and Neus Wang et al. (2021), implicitly represent signed distance functions (SDFs) and employ a SDF-density conversion function to encourage precise surface reconstruction. Other methods Li et al. (2023); Rosu & Behnke (2023); Wang et al. (2023); Yariv et al. (2023); Fu et al. (2022); Zhang et al. (2023) have also been proposed, introducing different representations and optimization techniques to further enhance reconstruction quality and efficiency. Nevertheless, these aforementioned methods struggle to handle indoor scenes with a large number of low-frequency areas (e.g., walls and floors), as photometric loss becomes unreliable in such regions.
@@ 3.1 Preliminaries
where $\alpha_{i}$ is the opacity of $i$-th ray segment, $T_{i}$ denotes the transmittance of light reaching the $i$-th segment, $\delta_{i}=t_{i}-t_{i-1}$ denotes distance between adjacent samples.
@@
$\sigma(\textbf{x})=\frac{1}{\beta}\Psi_{\beta}\left(-s(\textbf{x})\right)=\left\{\begin{array}[]{ll}\frac{1}{2\beta}\exp\left(\frac{-s(\textbf{x})}{\beta}\right)&\text{ if }s(\textbf{x})\leq 0\\ \frac{1}{\beta}-\frac{1}{2\beta}\exp\left(\frac{s(\textbf{x})}{\beta}\right)&\text{ if }s(\textbf{x})>0\end{array}\right.,$ (3)
@@
$\mathcal{L}_{\text{eik }}=\frac{1}{N}\sum_{i=1}^{N}\left(\left\|\nabla s\left({\textbf{x}}_{i}\right)\right\|_{2}-1\right)^{2},$ (6)
Following Neuralangelo Li et al. (2023), we further employ numerical gradients to enhance surface geometric consistency and use curvature loss for smoothness. The detailed definitions of them are provided in the appendix.
@@ 3.2 Normal deflection field
$\textbf{Q}(\textbf{r})=\sum_{i=1}^{N}T_{i}\alpha_{i}\textbf{q}_{i},$ (8)
@@
$\mathcal{L}^{d}_{{normal}}=\sum_{\textbf{r}\in\mathcal{R}}\|\hat{\textbf{N}}^{d}(\textbf{r})-\textbf{N}(\textbf{r})\|_{1}+\left\|1-\hat{\textbf{N}}^{d}(\textbf{r})^{T}\textbf{N}(\textbf{r})\right\|_{1},$ (10)
@@ 3.3 Adaptive deflection angle prior loss
As $\Delta\theta$ increases, $g^{d}(\Delta\theta)$ increases, and $g(\theta)$ decreases, since we should less apply priors in complex areas indicated by large deflection angles. The combined adaptive normal and depth prior loss is termed the adaptive deflection angle prior loss (as illustrated in Figure 2); see appendix for detailed definitions of the proposed modulation functions.
@@ 3.4 Deflection angle guided optimization
Textureless areas, such as walls, constitute a significant portion of indoor scenes. These regions converge rapidly by leveraging prior knowledge, partly due to their inherent simplicity. To prevent continuous oversampling of well-learned smooth regions, we proactively sample more rays in complex areas to capture finer details. The sampling process is naturally guided by the learned deflection angles (see supplementary for details).
@@
$\mathcal{L}^{d}_{color}=\sum_{\textbf{{r}}\in\mathcal{R}}{w_{color}({\Delta\theta}(\textbf{r}))\|\hat{\textbf{C}}(\textbf{r})-{\textbf{C}(\textbf{r})}\|_{1}},$ (14)
where ${\Delta\theta}(\textbf{r})$ denotes the actual deflected angle of the sampled ray r, and $w_{color}$ is the re-weighting function guided by ${\Delta\theta}$ (see appendix for details).
@@
where $\textbf{r}(t_{i})=\textbf{o}+t_{i}\textbf{v}$ represents a sampled point on the ray, and $f_{g}$ is the geometry network, which predicts the SDF value. This transformation effectively alleviates bias issues, as illustrated in Fig 3. However, we observe that simply applying it results in the overall surface failing to converge. Therefore, we only partially apply it to thin structures indicated by the learned deflection angles.
@@ 3.5 Optimization
$\mathcal{L}=\mathcal{L}_{color}^{d}+\lambda_{1}\mathcal{L}_{eik}+\lambda_{2}\mathcal{L}_{curv}+\lambda_{3}\mathcal{L}^{ad}_{normal}+\lambda_{4}\mathcal{L}^{ad}_{depth},$ (16)
@@ 4 Experiment
Figure 4: Qualitative results on ScanNet. Compared to previous state-of-the-art methods, our approach generates more thin and fine-grained structures, such as chair legs and detailed table layouts.
@@
MonoSDF Yu et al. (2022b) 3.17 3.70 13.75 5.68 6.58
Ours 4.37 6.45 16.28 10.30 9.35
Table 2: F-score on the TanksandTemples dataset, evaluated at a resolution of 768.
@@ 4.1 Comparisons
ScanNet: In Figure 4, we present qualitative results and in Table 1, quantitative results are displayed. These quantitative results are averaged over the selected four scenes. ND-SDF significantly surpasses previous best-performing works, achieving state-of-the-art performance. Specifically, we achieve the highest F-score, a reliable metric that reflects reconstruction accuracy by balancing both Accuracy and Completeness. When compared with HelixSurf and MonoSDF (as shown in Figure 4), our approach accurately captures thin structures, such as chair legs and lamp rings. The results highlight the effectiveness of our proposed deflection methods, which substantially enhance the surface accuracy and the recovery of thin and fine-grained structures in indoor environments.
@@
MonoSDF Yu et al. (2022b) 0.057 0.032 0.651 0.703 0.044 0.675
@@
MonoSDF Yu et al. (2022b) 92.11 2.94 86.18
Ours 92.85 2.57 91.6
Table 4: Quantitative results of Replica dataset, significantly surpassing previous methods.
Replica: Following MonoSDF Yu et al. (2022b), we used their pre-processed masks to filter out abnormal priors. Quantitative results are presented in Table 4, where our approach achieves the highest F-score and the lowest Chamfer Distance, significantly surpassing previous methods. Refer to the supplementary material for visualization results.
Tanks and temples: We also conducted experiments on the advanced split of T&T, which features several challenging large-scale indoor scenes. As shown in Table 2, ND-SDF surpasses previous methods that utilize priors in terms of quantitative performance. Additional qualitative results can be found in the supplementary material, where our method recovered substantial detailed structures.
@@
Table 5: F-score of Different Priors Models.
ScanNet++: ScanNet++ comprises numerous complex indoor scenes, each offering high-resolution GT mesh and RGB data. Since no public results are available for ScanNet++, we trained several baselines methods, including VolSDF, MonoSDF, and Baked-Angelo Yu et al. (2022a). Among these, only MonoSDF utilized monocular priors as we did. The quantitative results are summarized in Table 3, demonstrating the state-of-the-art performance of our method across all metrics. For additional visualization results, please refer to the appendix.
@@
Table 6: Quantitative Results of Ablation for Different Modules. ‘Base’ can be viewed as a simple combination of MonoSDF and Neuralangelo. ‘ND’ denotes applying the deflection field and the deflected normal loss term $L^{d}_{normal}$. ‘AP’ refers to the adaptive deflection angle prior loss terms including $L^{ad}_{normal}$ and $L^{ad}_{depth}$. ‘DO’ denotes the two deflection angle guided optimizations, namely sampling and photometric loss. ‘DU’ denotes the deflection angle guided unbiased rendering.
Figure 5: Visualization Results of the Models after Ablation. Detailed definitions of these models are provided in Table 6.
@@ A.3.3 Numerical Gradient
Following Neuralangelo Li et al. (2023), we employ numerical gradients w.r.t multi-resolution hash grids to promote surface consistency, as proved effective in their work. The numerical gradient is computed as follows (considering only the SDF value output of geometry network $f_{g}$):
@@
We also utilize the curvature loss Li et al. (2023) to further encourage the smoothness of the reconstructed surfaces. This loss term is defined as:
@@ B.2 Ablation of adaptive monocular depth loss
Method w/o $\mathcal{L}^{d}_{depth}$ w $\mathcal{L}^{d}_{depth}$
F-score 0.664 0.686
Figure 11: Visualization results on Replica dataset.

### v3

@@ 1 Introduction
3D surface reconstruction Chen et al. (2024a); Wang et al. (2024); Ye et al. (2022; 2023b); Liu et al. (2021); Li et al. (2020) aims to recover watertight, dense 3D geometry from multi-view images Hartley & Zisserman (2003), representing a significant research area in computer vision and graphics. The recovered surfaces have proven invaluable for many downstream tasks, including robotic navigation, AR/VR, and smart cities.
@@
However, recovering high-fidelity surfaces remains challenging, as relying solely on color images for supervision often results in an underconstrained problem, especially for regions like textureless walls and ceilings. Previous methods Yu et al. (2022b); Wang et al. (2022) have attempted to mitigate this issue by incorporating auxiliary supervision. For instance, MonoSDF employs monocular cues from a pretrained model as pseudo-ground truth to supervise the reconstruction process, which alleviates the issue in textureless regions. However, the large errors due to domain gaps in monocular cues and the inconsistencies in view-dependent prior guidance often lead to visible detail loss and erroneous surfaces Yu et al. (2022b).
@@ 2 Related Work
Recently, the representation of 3D scenes using neural fields has gained popularity due to their expressiveness and simplicity. NeRF-type approaches Mildenhall et al. (2021); Barron et al. (2022); Müller et al. (2022); Fridovich-Keil et al. (2022); Sun et al. (2022) have proposed encoding coordinate-based density and appearance of scenes by utilizing simple multilayer perceptrons (MLPs) and explicit structures such as voxel grids, resulting in photorealistic novel view synthesis. However, these approaches fail to accurately recover surfaces due to the lack of constraints on density. To address this issue, subsequent works, such as VolSDF Yariv et al. (2021) and Neus Wang et al. (2021), implicitly represent signed distance functions (SDFs) and employ a SDF-density conversion function to encourage precise surface reconstruction. Other methods Li et al. (2023b); Rosu & Behnke (2023); Wang et al. (2023); Yariv et al. (2023); Fu et al. (2022); Zhang et al. (2023) have also been proposed, introducing different representations and optimization techniques to further enhance reconstruction quality and efficiency. Nevertheless, these aforementioned methods struggle to handle indoor scenes with a large number of low-frequency areas (e.g., walls and floors), as photometric loss becomes unreliable in such regions.
@@ 3.1 Preliminaries
where $\alpha_{i}$ is the opacity of $i$-th ray segment, $T_{i}$ denotes the transmittance of light reaching the $i$-th ray segment, $\delta_{i}=t_{i}-t_{i-1}$ represents the distance between adjacent sample points.
@@
$\sigma(\textbf{x})=\frac{1}{\beta}\Psi_{\beta}\left(-s(\textbf{x})\right)=\left\{\begin{array}[]{ll}\frac{1}{2\beta}\exp\left(\frac{-s(\textbf{x})}{\beta}\right)&\text{ if }s(\textbf{x})\geq 0\\ \frac{1}{\beta}-\frac{1}{2\beta}\exp\left(\frac{s(\textbf{x})}{\beta}\right)&\text{ if }s(\textbf{x})<0\end{array}\right.,$ (3)
@@
$\mathcal{L}_{\text{eik }}=\frac{1}{N}\sum_{i=1}^{N}\left(\left\|\nabla s\left({\textbf{x}}_{i}\right)\right\|_{2}-1\right)^{2}.$ (6)
Following Neuralangelo Li et al. (2023b), we further employ numerical gradients to enhance surface geometric consistency and use curvature loss for smoothness. The detailed definitions of them are provided in the appendix.
@@ 3.2 Normal deflection field
$\textbf{Q}(\textbf{r})=\sum_{i=1}^{N}T_{i}\alpha_{i}\textbf{q}_{i}.$ (8)
@@
$\mathcal{L}^{d}_{{normal}}=\sum_{\textbf{r}\in\mathcal{R}}\|\hat{\textbf{N}}^{d}(\textbf{r})-\textbf{N}(\textbf{r})\|_{1}+\left\|1-\hat{\textbf{N}}^{d}(\textbf{r})^{T}\textbf{N}(\textbf{r})\right\|_{1}.$ (10)
@@ 3.3 Adaptive deflection angle prior loss
As $\Delta\theta$ increases, $g^{d}(\Delta\theta)$ increases, and $g(\Delta\theta)$ decreases, since we should less apply priors in complex areas indicated by large deflection angles. The combined adaptive normal and depth prior loss is termed the adaptive deflection angle prior loss (as illustrated in Figure 2); see appendix for detailed definitions of the proposed modulation functions.
@@ 3.4 Deflection angle guided optimization
Textureless areas, such as walls, constitute a significant portion of indoor scenes. These regions converge rapidly by leveraging prior knowledge, partly due to their inherent simplicity. To prevent continuous oversampling of well-learned smooth regions, we proactively sample more rays in complex areas to capture finer details. The sampling process is naturally guided by the learned deflection angles (see Appendix A.3.2 for details).
@@
$\mathcal{L}^{d}_{color}=\sum_{\textbf{{r}}\in\mathcal{R}}{w_{color}({\Delta\theta}(\textbf{r}))\|\hat{\textbf{C}}(\textbf{r})-{\textbf{C}(\textbf{r})}\|_{1}}.$ (14)
${\Delta\theta}(\textbf{r})$ denotes the actual deflected angle and $w_{color}$ is the re-weighting function (Appendix A.3.2).
@@
where $\textbf{r}(t_{i})=\textbf{o}+t_{i}\textbf{v}$ is a sampled point on the ray, and $f_{g}$ is the geometry network, which predicts the SDF value. This transformation effectively mitigates bias, as shown in Fig 3. However, we observe that simply applying it leads to poor convergence of the overall surface. Therefore, we only partially apply it to thin structures indicated by the learned deflection angles. A detailed analysis of the convergence issue is presented in Appendix B.1.
@@ 3.5 Optimization
$\mathcal{L}=\mathcal{L}_{color}^{d}+\lambda_{1}\mathcal{L}_{eik}+\lambda_{2}\mathcal{L}_{curv}+\lambda_{3}\mathcal{L}^{ad}_{normal}+\lambda_{4}\mathcal{L}^{ad}_{depth}.$ (16)
@@ 4 Experiment
Figure 4: Qualitative results on ScanNet. Our approach effectively generates thin structures, such as chair legs, and produces more accurate surfaces.
@@
MonoSDF (Grid) Yu et al. (2022b) 6.32 6.43 18.73 8.96 10.11
DebSDF (Grid) Xiao et al. (2024) 7.45 4.21 19.33 7.64 9.66
Ours 7.07 10.74 23.92 16.11 14.46
Table 2: F-score on the TanksandTemples dataset. ’Grid’ indicates using hash encoding as the feature representation backend. More detailed comparison results can be found in the appendix.
@@ 4.1 Comparisons
ScanNet: In Figure 4, we present qualitative results and in Table 1, quantitative results are displayed. These quantitative results are averaged over the selected four scenes, consistent with MonoSDF. ND-SDF significantly surpasses previous best-performing works, achieving state-of-the-art performance. Specifically, we achieve the highest F-score, a reliable metric that reflects reconstruction accuracy by balancing both Accuracy and Completeness. When compared with HelixSurf and DebSDF (as shown in Figure 4), our approach also accurately captures thin structures, such as chair legs. The results highlight the effectiveness of our proposed deflection methods, which substantially enhance the surface accuracy and the recovery of thin and fine-grained structures in indoor environments.
@@
MonoSDF (Grid) Yu et al. (2022b) 0.057 0.032 0.651 0.703 0.044 0.675
DebSDF (Grid) Xiao et al. (2024) 0.060 0.031 0.669 0.727 0.045 0.696
@@
MonoSDF (Grid) Yu et al. (2022b) 90.93 3.23 85.91
DebSDF (MLP) Yu et al. (2022b) 93.23 2.90 88.36
Ours 92.85 2.57 91.60
Table 4: Quantitative results of Replica dataset. Note that DebSDF reported results only with MLP representations.
Replica: Following MonoSDF Yu et al. (2022b), we used their pre-processed masks to filter out abnormal priors. Quantitative results are presented in Table 4, where our approach achieves the highest F-score and the lowest Chamfer Distance, significantly surpassing previous methods.
Tanks and temples: We also conducted experiments on the advanced split of T&T, featuring several challenging large-scale indoor scenes. As shown in Table 2, ND-SDF achieves the highest average F-score compared to previous methods. Additional comparison results can be found in Appendix C.3, where our method recovered substantial detailed structures.
@@
Table 5: F-score of Different Priors Models on ScanNet++.
ScanNet++: ScanNet++ comprises numerous complex indoor scenes, each offering high-resolution GT mesh and RGB data. Since no public results are available for ScanNet++, we trained several baselines methods, including VolSDF, MonoSDF, DebSDF and Baked-Angelo Yu et al. (2022a). Among these, MonoSDF and DebSDF utilized monocular priors as we did. The quantitative results, summarized in Table 3, demonstrate the state-of-the-art performance of our method, particularly in terms of the Recall metric. This highlights the capability of our approach to recover significantly finer structures of the GT surface. Additional visualization results are prensented in Appendix C.2.
@@
Table 6: Quantitative Results of Ablation for Different Modules on ScanNet++. ‘Base’ can be viewed as a simple combination of MonoSDF and Neuralangelo. ‘ND’ denotes applying the deflection field and the deflected normal loss term $L^{d}_{normal}$. ‘AP’ refers to the adaptive deflection angle prior loss terms including $L^{ad}_{normal}$ and $L^{ad}_{depth}$. ‘DO’ denotes the two deflection angle guided optimizations, namely sampling and photometric loss. ‘DU’ denotes the deflection angle guided unbiased rendering.
Figure 5: Visualization Results of the Models after Ablation on ScanNet++. Detailed definitions of these models and corresponding quantitative results are provided in Table 6.
@@ 5 Conclusion
Acknowledgments
This work was partially supported by Key R&D Program of Zhejiang Province (No.2023C01039) and NSF of China (No.62425209).
@@ A.3.3 Numerical Gradient
Following Neuralangelo Li et al. (2023b), we employ numerical gradients w.r.t multi-resolution hash grids to promote surface consistency, as proved effective in their work. The numerical gradient is computed as follows (considering only the SDF value output of geometry network $f_{g}$):
@@
We also utilize the curvature loss Li et al. (2023b) to further encourage the smoothness of the reconstructed surfaces. This loss term is defined as:
@@ B.1 Ablation of partial unbiased rendering
We further visualize the inverse variance ($1/\beta$) of the SDF-density conversion function (Eq. 3) during training in Figure 11(a) and find that, without partial unbiasing, $1/\beta$ fails to converge to a relatively high value. In fact, the larger the inverse variance, the sharper and more accurate the SDF-density field becomes. Conversely, a smaller value indicates greater uncertainty in the geometric surface, which leads to lower reconstruction accuracy and failure to converge properly. As shown in Figure 11(b), an increase in $1/\beta$ results in the rendering weight becoming more concentrated near the surface. Without a partial strategy, the surface is learned inaccurately and appears blurry.
More discussions about unbiasing methods are presented in Appendix C.4
Figure 11: (a) The inverse variance curve of the SDF-density conversion function during training. (b) The rendering weight curve with respect to the inverse variance.
@@
Method w/o $\mathcal{L}^{d}_{depth}$ w $\mathcal{L}^{d}_{depth}$
F-score 0.664 0.686
@@ B.2 Ablation of adaptive monocular depth loss
Figure 12: Detailed visualization comparisons of structural region confidence maps obtained using various methods. A comprehensive analysis is provided in Appendix B.3.
B.3 Confidence map analysis
As illustrated in Figure 12, we present the confidence maps that highlight structural regions, which are obtained using a variety of methods, including the NCC map Wang et al. (2022), edge map Li et al. (2023a), uncertainty map Xiao et al. (2024) and our angle map. The NCC map exhibits low performance due to its simplistic local cross-correlation computing method, and it fails to properly distinguish high-frequency flat regions, such as textured floors. The edge map, generated by a pretrained network Su et al. (2021), effectively captures object edges but still lacks precise localization of the fine structures. The uncertainty map, which is learned by constructing an uncertainty field, performs better than the first two methods. However, it only provides an approximate indication of high-frequency regions, as it inherently serves as an intensity field to measure the confidence of monocular cues, independent of actual geometry. As a result, the uncertainty map fails to precisely evaluate the actual differences between scene geometry and priors, leading to inefficient utilization of various priors. Additionally, it fails to recognize highlighted areas, as shown in the second row of Figure 12 and the first row of Figure 16. In contrast, our method, by modeling geometry-relevant SO(3) residual fields, accurately locates specific structural regions and provides precise angular deviations. The angle map clearly reflects the underlying geometry difference. Compared to the methods discussed above, our approach yields more reasonable and accurate results.
In terms of the sampling strategy guided by the confidence maps illustrated in Figure 12, our sampling process is fully weighted by geometry-aware angular deviations and is precisely localized to structural regions. In contrast, other methods, such as edge-aware and uncertainty-driven sampling, exhibit lower localization precision and are less sensitive to geometric differences from the priors, failing to provide effective sampling weights for various priors.
This analysis is further validated through extensive experiments. As shown in Figure 16 and Figure 17, our method generates more accurate fine structures compared to DebSDF.
B.4 Analysis of the Robustness of Monocular Cues
To deeply investigate the effectiveness of our proposed explicit deflection-based method, we conducted experiments under the assumption that the priors for smooth regions are inaccurate. Specifically, we manually rotated all monocular normals in the smooth regions by angles ranging from 0° to 60°. A SAM model Medeiros et al. (2024) was employed to detect the flat areas. The rotation was performed by rotating the normal priors by a specified degree (e.g., 60°) toward a pre-defined axis. We evaluated MonoSDF, DebSDF, and ND-SDF on the 0e75f3c4d scene from ScanNet++. The results are summarized in Table 10 and Figure 13. As demonstrated in the results, decreases in F-score are observed across all methods. Specifically, MonoSDF produces an entirely abnormal result, as it rigidly relies on the biased priors. DebSDF also fails to maintain robustness in this scenario. Although the uncertainty field in DebSDF can easily mask irregular priors by learning uncertainty variances, it leads to inefficient utilization of the consistency information inherent in monocular cues within smooth regions. In contrast, our method explicitly models the actual bias amplitude, achieving notably robust performance with the smallest decrease in F-score (Table 10), even when the normal priors are biased by as much as 60°. These findings further validate the generalization and robustness of our method in non-trivial assumptions of the monocular cues.
Figure 13: Visualization results with monocular normal cues manually biased by 60 degrees. Our method accurately captures the view-dependent biased priors and yields promising reconstruction result.
Methods 15° 30° 45° 60° $\rightarrow$60°
Ours 0.750 0.735 0.678 0.654 0.709 0.655 -0.095
MonoSDF 0.633 / / / / 0.140 -0.493
DebSDF 0.701 / / / / 0.517 -0.184
Table 10: Quantitative results under manually biased monocular normal cues. The observed minimal decline in F-score underscores the effectiveness and robustness of our approach.
Appendix C More results
C.1 Replica
We present additional visualization results of our method on Replica in Figure 14.
Figure 14: Qualitative results of our method on Replica dataset.
C.2 ScanNet++
We present detailed quantitative results of the selected six scenes from ScanNet++ in Table 11 and provide more visual comparison results with previously state-of-the-art methods in Figure 15, Figure 16 and Figure 17. Compared to the baselines, our method yields significantly improved reconstructions both visually and quantitatively, notably excelling at capturing thin, fine-grained structures while maintaining smoothness in flat regions.
C.3 Tanks and temples
As illustrated in Figure 19 and Figure 20, we extracted the reconstructed meshes at a resolution of 2048. At this high resolution, the meshes demonstrate that our method significantly enhances the generation of small and thin structures, even in challenging large-scale indoor scenes. For all experiments, the hash grid settings were kept at their defaults, with resolutions ranging from $2^{5}$ to $2^{11}$ across 16 levels, and a hashmap size of $2^{19}$.
We present more qualitative comparisons with previous state-of-the-art baselines, including MonoSDF Yu et al. (2022b) and DebSDF Xiao et al. (2024) at 2048 resolution, as depicted in Figure 18. It is evident that ND-SDF achieves accurate dense surfaces replete with extensive details.